a) A wire of a certain length (a=0.0065/°C) has a resistance of 4Ω (at 20°C). Find its resistance at 80°C.



1
Expert's answer
2021-11-01T19:24:44-0400

"R=R_0(1+aΔT) \\
R = 4(1 + 0.0065 \times (80-20)) \\
R = 4 \times 1.39 = 5.56 \; \Omega"